I don't know where to put this... so I just tried here... Don't know how to do this promble at all... I'd love it if someone told me how..
1/3-6v=2 2/3

Answer by checkley71(8403)

HOPE THIS EXPLANATION IS ADEQUATE FOR YOU TO UNDERSTAND HOW TO SOLVE THIS PROBLEM.
IF NOT, THE LET ME KNOW.
-------------------------------------------------------------------------
1/3-6v=2 2/3 now subtract -1/3 from both sides thus eliminating the 1/3 fraction from the left side.
-6v=2 2/3-1/3 now combine (subtract 1/3 from 2 2/3)
-6v=2 1/3 convert the mixed fraction to an improper fraction.
-6v=7/3 divide both sides by -6 to isolate the v term.
v=(7/3)/-6 dividing a fraction by an integer requires you to invert the integer [-6 becomes -1/6) & multipy.
v=(7/3)(-1/6) now multiply both numerators & denominators.
v=-7/18 answer.
proof:
1/3-6(-7/18=2 2/3
1/3+42/18=2 2/3
1/3+2 1/3=2 2/3
2 2/3=2 2/3